A researcher is interested in seeing if the amount of protein in a pigeon's feed influences the rate at which the pigeon pecks a key to receive the feed as a reward. She makes up three different types of feed: Type A, which has a large amount of protein; Type B, which has a moderate amount of protein; and Type C, which has very little protein. The researcher hypothesizes that pigeons will peck faster for the low protein feed because they need more of it to meet their nutritional requirements.        

Type A

Type B

Type C

20

30

20

25

35

35

25

35

40

20

35

40

30

35

45



** The SSbg (sums of squares between groups) is 413.333. The SSwg (sums of squares within groups) is 460.000.

The degrees of freedom within groups equal ______.

Select one:

a. 12

b. 14

c. 4

d. 3

e. 2
Answer #1

The degrees of freedom within groups equal 12

Option a) is correct.

within groups degree of freedom = n - k = (5*3) - 3 = 12